How do you write the equation y=2/3x -2 in standard form? I think it is -2/3x + y = -2
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

y=2/3x -2


3y=2x - 6 ... Note: Multiply EVERY term by the LCD 3 to clear out the fraction


-2x + 3y = -6


2x - 3y = 6


So the answer is 2x - 3y = 6

Answer by stanbon(75887)

How do you write the equation y=2/3x -2 in standard form?
I think it is -2/3x + y = -2
---
Standard Form: ax + by = c where a,b, and c are integers
---
y=2/3x - 2
Multiply thru by 3 to get:
3y = 2x - 6
---
Rearrange:
2x - 3y = 6
===================
Cheers,
Stan H.